# BCQ: Block Clustered Quantization for 4-bit (W4A4) LLM Inference

### v1

@@ 4.2.2 Accuracy on LM evaluation harness tasks
Across 00-shot LM evaluation harness tasks in Table 4, LO-BCQ shows significant improvement in average accuracy compared to MX, MXFP and VSQ at equivalent bitwidth. Further, across models during $4.5$-bit quantization, LO-BCQ achieves $<1$% loss in average accuracy compared to the respective unquantized baselines. When the bitwidth of LO-BCQ is increased by varying its configuration, we find that the average accuracy generally increases albeit with a few exceptions. Although these variations are small ($<0.5$%), we believe that they arise due to the universal calibration of codebooks. Our codebooks are calibrated on a batch of training data from the Wikitext-103 dataset and the GPT3-126M model and remain frozen across all datasets and models.
Although the focus of this work is weight+activation quantization, we also compare to prior art weight-only quantization proposals for completeness. Table 3 compares weight-only (W4A8) LO-BCQ with a block array size of $128$ to other weight-only quantization proposals such as GPTQ and AWQ of comparable block array size and effective bit-width. As shown, LO-BCQ with $2$, $4$, $8$ and $16$ codebooks with effective bitwidth of $4.19$, $4.31$, $4.44$ and $4.56$, respectively, achieves significantly lower perplexity loss. It is worth noting that we evaluate this loss on Wikitext-103 dataset, which is a much larger dataset compared to Wikitext2 used by other works. Further, we compare LO-BCQ against other codebook-based quantization methods such as Quip# and AQLM. As shown, LO-BCQ achieves comparable accuracy using only a small codebook size ($8$ codebooks with $16$ entries each) compared to significantly larger codebook sizes ($2^{16}$ codebooks with $8$ entries each) required by QuiP# and AQLM.
@@ 4.2.3 Accuracy on MMLU tasks
Similarly, in $5$-shot MMLU tasks LO-BCQ achieves $<1$% loss in average accuracy with $4.5$-bits per scalar compared to respective unquantized baselines across GPT3-22B and Llama2-70B models. Further, LO-BCQ achieves a significantly better accuracy compared to all of our block quantization baselines such as VSQ, MX and MXFP4 at equivalent bitwidth. Across Llama2 models, LO-BCQ with a smaller bitwidth ($4.25$-bits) outperforms VSQ and MX4 with a comparatively larger bitwidth ($4.5$-bits). While the $0.5$-bit overhead in VSQ and MX4 are used on per-block array scale-factors, the $0.25$-bit overhead of LO-BCQ is shared between scale-factors and codebook selectors. Therefore, the superior accuracy of LO-BCQ can be attributed to the better representation by selecting the best codebook for each block.
@@ 4.3 Ablation Studies
Table 6 shows the perplexity of LO-BCQ on Wikitext-103 dataset and across Llama2-70B and GPT3-22B models when its configuration is varied. For a given $L_{b}$ (block length), larger number of codebooks results in better perplexity. This is intuitive since larger number of codebooks leads to better representation of the values in each block since LO-BCQ allows it to map to the codebook with best representation. Further, when the block array size is reduced, we achieve better perplexity. The block array corresponds to the granularity of normalization. As discussed in section 2.3, normalization improves convergence of LO-BCQ and results in better perplexity. When comparing configurations with same bitwidth, we find that the configuration with larger number of codebooks is better than smaller block array. This shows that the per-block metadata is better utilized for codebook selectors than scale factors. Furthermore, we find that reducing the block length ($L_{b}$) below $8$ results in diminishing returns. This is because, the overhead of storing codebook selectors is larger for a smaller block. For a given bitwidth, configuration with smaller $L_{b}$ has fewer codebooks. Therefore, these configurations result in larger loss in perplexity.
@@ A.1 Lloyd-Max Algorithm
Figure 7 compares the quantization levels of a $7$-bit floating point (E3M3) quantizer (left) to a $7$-bit Lloyd-Max quantizer (right) when quantizing a layer of weights from the GPT3-126M model at a per-tensor granularity. As shown, the Lloyd-Max quantizer achieves substantially lower quantization MSE. Further, Table 8 shows the superior perplexity achieved by Lloyd-Max quantizers for bitwidths of $7$, $6$ and $5$. The difference between the quantizers is clear at 5 bits, where per-tensor FP quantization incurs a drastic and unacceptable increase in perplexity, while Lloyd-Max quantization incurs a much smaller increase. Nevertheless, we note that even the optimal Lloyd-Max quantizer incurs a notable ($\sim 1.5$) increase in perplexity due to the coarse granularity of quantization.
@@ A.4.3 Quantization Scheme
In the above equation, ||$\cdot$|| ⋅ | denotes the absolute value function.

### v2

@@ 4.2.2 Accuracy on LM evaluation harness tasks
Across 0-shot LM evaluation harness tasks in Table 6, LO-BCQ shows significant improvement in average accuracy compared to MX, MXFP and VSQ at equivalent bitwidth. Further, across models during $4.5$-bit quantization, LO-BCQ achieves <1<1% loss in average accuracy compared to the respective unquantized baselines. When the bitwidth of LO-BCQ is increased by varying its configuration, we find that the average accuracy generally increases albeit with a few exceptions. Although these variations are small (<0.5<0.5%), we believe that they arise due to the universal calibration of codebooks. Our codebooks are calibrated on a batch of training data from the Wikitext-103 dataset and the GPT3-126M model and remain frozen across all datasets and models.
Although the focus of this work is weight+activation quantization, we also compare to prior art weight-only quantization proposals for completeness. Table 4 compares weight-only (W4A8) LO-BCQ with a block array size of $128$ to other weight-only quantization proposals such as GPTQ and AWQ of comparable block array size and effective bit-width. As shown, LO-BCQ with $2$, 44, 88 and $16$ codebooks with effective bitwidth of $4.19$, $4.31$, $4.44$ and $4.56$, respectively, achieves significantly lower perplexity loss. It is worth noting that we evaluate this loss on Wikitext-103 dataset, which is a much larger dataset compared to Wikitext2 used by other works. Further, we compare LO-BCQ against other codebook-based quantization methods such as Quip# and AQLM. As shown, LO-BCQ achieves comparable accuracy using only a small codebook size ($8$ codebooks with $16$ entries each) compared to significantly larger codebook sizes ($2^{16}$ codebooks with $8$ entries each) required by QuiP# and AQLM.
Table 5 shows the Wiki2 perplexity results on weight-only quantization of Llama2 models with LO-BCQ and compares against state-of-the-art 2-bit and 3-bit quantization proposals such as QuIP# and AQLM. We find that despite the significantly smaller number of codebooks (<=8<=8 with 16 entries each) utilized by LO-BCQ compared to about $2^{12}$ to $2^{16}$ codebooks with 8 entries each in the other methods, LO-BCQ achieves competitive results. During 3-bit quantization, LO-BCQ with effective bitwidth of 3.375 bits achieves lower perplexity than QuIP# where LDLQ Tseng et al. (2024) is applied to both methods and no finetuning is performed. Similarly, when compared against 2-bit QuIP#, we find that LO-BCQ with effective bitwidth of 2.375 bits achieves better Wiki2 perplexity. Further, 3.5-bit and 2.5-bit LO-BCQ without finetuning suffers only a small loss compared to finetuned AQLM during 3-bit and 2-bit quantization of weights, respectively.
@@ 4.2.3 Accuracy on MMLU tasks
Similarly, in $5$-shot MMLU tasks LO-BCQ achieves <1<1% loss in average accuracy with $4.5$-bits per scalar compared to respective unquantized baselines across GPT3-22B and Llama2-70B models. Further, LO-BCQ achieves a significantly better accuracy compared to all of our block quantization baselines such as VSQ, MX and MXFP4 at equivalent bitwidth. Across Llama2 models, LO-BCQ with a smaller bitwidth ($4.25$-bits) outperforms VSQ and MX4 with a comparatively larger bitwidth ($4.5$-bits). While the $0.5$-bit overhead in VSQ and MX4 are used on per-block array scale-factors, the $0.25$-bit overhead of LO-BCQ is shared between scale-factors and codebook selectors. Therefore, the superior accuracy of LO-BCQ can be attributed to the better representation by selecting the best codebook for each block.
@@ 4.3 Ablation Studies
Table 8 shows the perplexity of LO-BCQ on Wikitext-103 dataset and across Llama2-70B and GPT3-22B models when its configuration is varied. For a given $L_{b}$ (block length), larger number of codebooks results in better perplexity. This is intuitive since larger number of codebooks leads to better representation of the values in each block since LO-BCQ allows it to map to the codebook with best representation. Further, when the block array size is reduced, we achieve better perplexity. The block array corresponds to the granularity of normalization. As discussed in section 2.3, normalization improves convergence of LO-BCQ and results in better perplexity. When comparing configurations with same bitwidth, we find that the configuration with larger number of codebooks is better than smaller block array. This shows that the per-block metadata is better utilized for codebook selectors than scale factors. Furthermore, we find that reducing the block length ($L_{b}$) below $8$ results in diminishing returns. This is because, the overhead of storing codebook selectors is larger for a smaller block. For a given bitwidth, configuration with smaller $L_{b}$ has fewer codebooks. Therefore, these configurations result in larger loss in perplexity. Table 10 compares perplexity achieved by LO-BCQ codebooks with INT4, INT6 and INT8 entries on Llama2-7B model and Wikitext-103 datatset. As shown, LO-BCQ with INT6 achieves negligible perplexity degradation compared to INT8, while INT4 suffers significantly larger degradation. As a result, we quantize LO-BCQ codebooks to INT6.
Table 10: Quantizing LO-BCQ codebook entries to INT4 vs INT6 vs INT8 on Wikitext-103. Perplexity is measured on the Llama2-7B model (BF16 PPL = 5.06).
.   Method Bitwidth of codewords INT4 INT6 INT8 LO-BCQ (g128, $N_{c}=2$) 6.24 5.38 5.35 LO-BCQ (g128, $N_{c}=4$) 6.21 5.27 5.25 LO-BCQ (g128, $N_{c}=8$) 6.20 5.21 5.21 LO-BCQ (g128, $N_{c}=16$) 6.18 5.21 5.19
@@ A.1 Lloyd-Max Algorithm
Figure 8 compares the quantization levels of a $7$-bit floating point (E3M3) quantizer (left) to a $7$-bit Lloyd-Max quantizer (right) when quantizing a layer of weights from the GPT3-126M model at a per-tensor granularity. As shown, the Lloyd-Max quantizer achieves substantially lower quantization MSE. Further, Table 11 shows the superior perplexity achieved by Lloyd-Max quantizers for bitwidths of $7$, 66 and $5$. The difference between the quantizers is clear at 5 bits, where per-tensor FP quantization incurs a drastic and unacceptable increase in perplexity, while Lloyd-Max quantization incurs a much smaller increase. Nevertheless, we note that even the optimal Lloyd-Max quantizer incurs a notable ($\sim 1.5$) increase in perplexity due to the coarse granularity of quantization.
@@ A.4.3 Quantization Scheme
In the above equation, $|\cdot|$ denotes the absolute value function.